OpenStudy (anonymous):

evaluate cot(cos^-1(-15/17))

OpenStudy (anonymous):

Enter your answer as a fraction using the slash bar ( / ).

OpenStudy (tkhunny):

Have you considered drawing a right triangle?

OpenStudy (anonymous):

i got -15/8 but not sure if correct or not

OpenStudy (anonymous):

Draw a right angled triangle by letting \[\theta=\cos^{-1}-\frac{15}{17}\] 

OpenStudy (anonymous):

then find the missing side, so that you can find cot.

OpenStudy (anonymous):

cot(theta)*

OpenStudy (anonymous):

missing side is 16 ?

OpenStudy (anonymous):

Use Pythagoras: \[a^2+b^2=c^2\]

OpenStudy (anonymous):

i think i can get it now thanks :)

OpenStudy (anonymous):

No worries. And to determine whether you get a positive or negative value, use your knowledge of ASTC.